Indiana University Multi-campus Modular Computing Curriculum

Introduction to Computing

Module # 1: Free Speech on the Internet

Prerequisite Skills

• Basic understanding of Internet and World Wide Web.

Goals

• Understand ethical and legal status of speech on the Internet.

1 — An Interview with Judge No Yt-Awl by Rennie the Mouse as Published in ChronOS Magazine, June 31, 1997—

Rennie the Mouse: Last week, the Supreme Court decided that the Communications Decency Act of 1996 (CDA) was unconstitutional. What was the CDA, and in what way did it violate the constitution?

Judge No Yt-Ol: The CDA was a part of the Telecommunications Act of 1996, which was signed into law by President Clinton on February 8, 1996. Two provisions of the Act were intended to protect minors from obscene, indecent, and offensive materials on the Internet.

Section 223(a) (the “indecent transmission” provision) makes it a crime to knowingly transmit, “by means of a telecommunications device,” any message or image “which is obscene or indecent,” to any person under the age of 18. (The term “telecommunications device” includes computer communications devices, such as modems, so the law applies to the Internet.)

Section 223(d) (the “patently offensive display” provision) makes it a crime to knowingly send or display, using an “interactive computer service,” any message or image

Module # 1: Free Speech on the Internet

1-1 which, “in context, depicts or describes, in terms patently offensive as measured by contemporary community standards, sexual or excretory activities or organs,” to any person under the age of 18.

Both sections provide for fines and/or imprisonment for up to two years.

The Supreme Court held that the CDA’s criminalization of indecent and offensive communication via the Internet unconstitutionally violates the freedom of speech protected by the First Amendment. However, insofar as the Act prohibits obscenity, it does not violate free speech. The First Amendment does not protect obscenity. Obscenity and child pornography were illegal even before the CDA, and they still are.

Mouse: I didn’t know there was a distinction between indecency and obscenity. What’s the difference?

Judge No: Obscenity is a more narrowly defined term. The Supreme Court in Miller v. California (1973) established the following definition of obscenity. A work is obscene if:

1. “the average person, applying contemporary community standards would find that the work, taken as a whole, appeals to the prurient interest [in sex];” (prurient means arousing lustful, lewd, lascivious thoughts or desires.)

2. “the work depicts or describes, in a patently offensive way, sexual conduct specifically defined by the applicable state law;” and

3. “the work, taken as a whole, lacks serious literary, artistic, political, or scientific value.”

The CDA doesn’t define “indecent” at all, but presumably there are a lot of materials that could be considered indecent but are not obscene.

Mouse: What are some materials that might be indecent or offensive but not obscene?

Judge No: Two of the twenty plaintiffs who challenged the CDA were Stop Prisoner Rape and Critical Path AIDS Project. Stop Prisoner Rape posts information, including graphics and statistics, concerning the occurrence and prevention of rape in prisons. Critical Path AIDS Project provides information on “safer sex” and how HIV is transmitted; it uses crude language to be sure that teenagers understand it.

Mouse: Besides obscenity, are there any other forms of unprotected speech?

Judge No: Certainly. Libel and slander—telling lies about someone that injure her reputation— are unprotected speech. Fraud is unprotected. Inciting a riot is unprotected speech. There is no right to cry, “Fire!” in a crowded theater where there is no fire.

Mouse: Why isn’t the government allowed to restrict indecency on the Internet, since it’s allowed to do so on radio and television?

Judge No: Why is the government allowed to regulate radio and TV at all? The answer lies in the nature of the medium. With broadcast radio and TV, there is a limited number of channels. If two stations in the same area broadcast on the same channel, their signals would interfere with each other. In this respect, the broadcast media are significantly different from the printed or spoken

1-2 word, and that gives government some justification both for licensing channels to specific broadcasters and, even after assigning a license, regulating the content of broadcasts in limited ways. For example, the government may require a station to offer educational or news programming, and it may restrict indecent speech to hours when children are not listening.

Another factor which makes radio and TV subject to regulation is their pervasive influence in American homes. Any child can turn on a TV set, even before learning to read. If TV were not regulated, a young child could turn on the tube and accidentally see some shocking, indecent thing.

Neither of these conditions, which make broadcast media the least protected by the First Amendment, applies to the Internet. First, there is no scarcity of Internet sites. One person’s setting up an Internet site does not in any way interfere with another’s. Second, it requires a little more sophistication to operate as a listener on the Internet. Even using the World-Wide Web, which is the easiest tool there is, generally requires the ability to read. So it is unlikely that a very young child is going to accidentally pick up some offensive or indecent material that way.

Mouse: What about older children? Don’t we need to protect them from Internet pornography?

Judge No: We certainly do, and the government, as well as parents, have a powerful interest in this. But—and here we are getting at the heart of the Court’s decision—although government has a compelling interest in protecting minors, the means it chooses must be “narrowly tailored” to that end, when it comes to First Amendment issues. In other words, in protecting youths, we must not chill the legitimate free expression of adults to other adults. We must not bring the level of adult communication down to the level of children.

The problem is that there is no practical means for an Internet content provider (speaker, publisher) to filter its material so that it is guaranteed to be readable only by adults. For most means of Internet communication—email sent to lists, chat, news, FTP, etc.—it is just technologically impossible. For one means of Internet communication—the World-Wide Web—it is technologically possible, by screening the offensive materials with a form that asks for, say, a credit card number or adult ID and then runs a program to verify it.

Ironically, most commercial pornographers on the Internet already use credit card or adult ID verification, so they would be unaffected by the CDA! But such techniques are too costly for many non-profit content providers (for example, Stop Prisoner Rape, Critical Path AIDS Project, the Carnegie Library). Faced with the choice between a possible criminal prosecution and prohibitively expensive age verification, many content providers would be forced to become silent.

This would be a terrible thing, because the Internet offers the most participatory form of communication the world has ever seen.

Judge Dalzell’s opinion in the District Court makes this clear. Dalzell pointed out four characteristics of the Internet which make it deserve the highest constitutional protection. Let me paraphrase Judge Dalzell’s remarks. First, it is very easy to get onto the Internet. Second, it is as easy to become a speaker on the Internet as it is to become a listener.

You have to be a millionaire to become a newspaper publisher, but with only a few thousand dollars, you can become an Internet publisher. This is an important point, because the barriers to entry in the print and broadcast media have led to a concentration of power, the power to shape and inform public opinion, into very few hands. Economic power has become almost equivalent

1-3 to influence.

Third, because it is so easy for ordinary people to become speakers on the Internet, the Internet provides an amazingly rich diversity of content. Fourth, all who wish to speak on the Internet have significant and approximately equal access. A factory worker can speak as much, and almost as well, as an industrial tycoon.

On the other hand, parents have options available to them. First, there are a number of childprotection software packages available, and the market demand for these means that they are going to get better and more numerous.

Examples include CyberPatrol, NetNanny, ParentalGuidance, SurfWatch, Netscape Proxy Server, The Internet Filter, and CYBERsitter. These programs can be set up at home to block access to selected sites or categories of sites, or to permit access to only selected sites. They are not perfect—for example, it is not currently possible for software to recognize and block sexually explicit pictures that have no accompanying suggestive text.

Nevertheless, they do provide a reasonably effective method for parents to safeguard their children. The second option, of course, is direct supervision of children on the Internet. Or don’t let them use it until they reach a suitable age.

For all these reasons—the absence of effective means for Internet content providers to limit the availability of their materials, the availability of such means for parents, the non-invasive nature of Internet communication, and the wonderful accessibility of the Internet for such a large and diverse group of speakers—the Supreme Court held that the Internet deserves the highest form of constitutional protection for speech.

Mouse: Does this mean that the government can do nothing at all?

Judge No: The government cannot limit indecency or offensive speech on the Internet, but there’s plenty that it can do. It can vigorously enforce the existing laws against obscenity and child pornography, which apply as much to the Internet as to any other medium of communication. It can also help to educate the public about the dangers and benefits of the Internet.

Mouse: Now that the offending portions of the CDA are dead, can we relax?

Judge No: Absolutely not. Some of the states have passed laws very similar to the CDA. The Congress may try something similar to CDA in the future. As Thomas Jefferson said: “Eternal vigilance is the price of freedom.”

2 Comments

1. The “patently offensive display” provision of the CDA defined “offensive” in a way that was limited to matters “sexual or excretory.”

Consequently, it does not apply to other forms of offensive speech, such as hate messages directed against racial, ethnic, or religious groups; recruitment for exploitative religious cults; or advice for terrorists, such as how to make bombs.

Such messages, are, obviously, morally repulsive, whether posted on the Internet or elsewhere. Equally obviously, they cannot be prohibited by the government under our constitutional system which guarantees freedom of speech. The best way to combat bad ideas is

1-4 with better ideas.

Incidentally, it is arguable that pornography—some or all of it—is itself a form of hate speech against women. (See, e.g., Fedler 1995)

2. Judge No’s statement that the government can vigorously enforce existing laws against obscenity echoes Judge Sloviter’s opinion in the District Court decision about the CDA.

In fact, while Congress was considering the CDA, the Justice Department itself expressed to Congress its view that the CDA was not necessary because existing laws against obscenity and child pornography gave it ample weapons for prosection.

Nevertheless, the three-pronged test of obscenity defined in Miller v. California has made it very difficult to prosecute cases for obscenity, so that prosecutions are rare, and convictions are even rarer. (Schauer 1995)

3. The global presence of the Internet raises another problem in connection with the “community standards” referred to in the Miller definition: what is the relevant community? In U. S. v. Thomas and Thomas, a United States Court of Appeals held that obscene materials sent from Milpitas, California to Memphis, Tennessee may be prosecuted “in any district from, through, or into which” the materials are sent, and that relevant community standards are those of “the community where the trial takes place” (in this case, Memphis).

The Thomases, operating a computer bulletin board in California, could have decided not to send materials to Tennessee, and so not be subject to Tennessee’s more stringent standards. However, the decision not to send material to a certain geographic region, any more than to a certain age group, is typically not within the power of content providers on the World-Wide Web and most other Internet services, so it is doubtful whether this decision would apply to the Internet.

On the other hand, perhaps commercial pornographers on the Internet, using credit card verification and adult ID, would be able to limit their distribution geographically as well as by age.

3 References

American Civil Liberties Union v. Reno, and American Library Association v. United States Department of Justice

1

(U. S. District Court decision about CDA, 1996),

Fedler, Joanne (1995). \underline{“A Feminist Critique of Pornography.”}^{2}

Miller v. California (1973). 413 U.S. 15. (Supreme Court case defining the current test for obscenity)

\underline{Reno v. American Civil Liberties Union}^{3} (Supreme Court decision about CDA, 1997),

Schauer, Frederick (1995). \underline{“The American Approach to the Law of Obscenity.”}^{4}

1-5

1http://www.aclu.org/court/cdadec.html 2http://wn.apc.org/fxi/books/chap2.htm 3http://www.aclu.org/court/renovacludec.html 4http://wn.apc.org/fxi/books/chap3.htm

\underline{U. S. v. Thomas and Thomas (1996).}^{5}^{5}http://www.aclu.org/court/thomas.html

1-6